OpenStudy (anonymous):

A bag has 2 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a green marble, do not replace it, and pick another green marble? A. 3/28 B. 1/28 C. 9/56 D. 1/8

OpenStudy (anonymous):

what is the probability you pick a green marble first?

OpenStudy (anonymous):

2/8 or 1/4

OpenStudy (anonymous):

ok good now that marble is gone what is the probability that the second one is green given that one green marble was chosen and not replaced?

OpenStudy (anonymous):

1/7?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

final answer you get via \[\frac{1}{4}\times \frac{1}{7}\]